From plucinsk@nesvig.harvard.edu Sun May 11 18:20:15 1997 Received: from space.mit.edu by wiwaxia AA17539; Sun, 11 May 97 18:20:13 EDT Received: from head-cfa (head-cfa.harvard.edu) by space.mit.edu AA02585; Sun, 11 May 97 18:20:06 EDT Received: from nesvig by head-cfa (SMI-8.6/SMI-SVR4) id SAA00311; Sun, 11 May 1997 18:18:55 -0400 Received: by nesvig (SMI-8.6/SMI-SVR4) id SAA29932; Sun, 11 May 1997 18:20:09 -0400 Date: Sun, 11 May 1997 18:20:09 -0400 From: plucinsk@nesvig.harvard.edu (Paul Plucinsky) Message-Id: <199705112220.SAA29932@nesvig> To: asc@nesvig.harvard.edu Subject: XRCF update Status: R XRCF Update ACIS Flat field measurements: Measurements are going smoothly during this phase. The ACIS instrument and the facility are working well. The pace is much more relaxed than previous phases since the measurements are typically longer and there are not many people here. We have completed the measurements of all ten CCDs at Si-K, Fe-K, Fe-L alpha. We have collected about half of the desired data at O-K and will complete those measurements in a couple of days. We are currently acquiring Cu-K data. The official stop time is late thursday but it appears that we will extend until saturday 6am. Some highlights so far: - measurements were made at -60 C as ACIS was cooling down to provide a baseline for the planned stray light measurements at TRW - Jim Francis ran ACIS through a mode survey to demonstrate that the instrument does indeed work in all planned modes and to produce telemetry in these modes. Unfortunately, the DLRS/MCC were not accepting data during the beginning of this test. - the ACIS internal calibration source was observed for an extended period - Steve Jones presented some preliminary count rate results showing that the FI CCDs agree with each other to within 1.5% at Si-K if one assumes the same QE for all chips. Once again demonstrating how uniform the FI CCDs are. The differences are larger at O-K where we are more sensitive to differences in the gate structure and the filter transmission. The BI CCDs don't agree so well with the FI CCDs assuming the same QE for both BI CCDs. But the BI CCDs are showing ~3.5 times higher detection efficiency than the FI CCDs as expected. - Jon Woo showed the average bias frames computed by the flight SW as the instrument was cooling down. We are investigating possible new hot pixels in the S1 BI CCD. We will need to compare data from lab measurements, the LL thermal vac test, and the ACIS+HRMA XRCF tests. - Catherine Grant and George Chartas showed that we are detecting Mn K X-rays at the top of the I0 and I1 CCDs with the door open, presumably from the internal source. We have a 12 hour background measurement coming up in which we should be able to measure this more accurately. - After a sequence of bizarre commands to the instrument, a bug in the flight SW may have been detected. During two separate runs, two different CCDs starting producing a fixed number of counts per frame with pulse heights below the event threshold. A re-boot cleared the problem. Jim Francis is thinking hard about this. We have been working primarily on the second floor (Kenny, Kester, myself, Steven Buczkowski is here and runs his first shift tomorrow, Norbert arrives tomorrow) analyzing the high speed tap data as quickly as possible. Jon has been the sole member of the 3rd floor analysis team working with the telemetry data. The test conductors are rejoicing that they can go home earlier than expected since the HRC team has decided not to do their flat field measurements here at the XRCF. 6 more days for ACIS, Paul